OpenStudy (anonymous):

can anyone help me integrate x square root(1-x^2)\ by substitution

OpenStudy (anonymous):

i think this is a straight forward u-sub

OpenStudy (anonymous):

\[u=1-x^2\] \[du=-2xdx\] \[-\frac{1}{2}\int \sqrt{u}du\] etc

OpenStudy (anonymous):

but what about the x